OpenStudy (anonymous):

Suppose |A| = 0, and that x1 is a particular solution of the system Ax = B. Show that any other solution x2 of this system can be written as x2 = x1 + x0, where x0 is a solution of the system Ax = 0

OpenStudy (anonymous):

Ax1 = B Ax2 = B subtract these two. Ax2-Ax1=B-B Take A as common A(x2-x1)=0 But its given that Ax0=0 Therefore equating, we have x2-x1 = x0 Therefore x2 = x1 + x0. I think i might be missing some subtle point in this proof. I hope somebody goes through this solution and comments.